OpenStudy (anonymous):

Solve the following problem using substitution or elimination: To skate at Roller Heaven, each person must pay a monthly membership fee (m) and a fee for each session (f). In January, Art attended 5 sessions and paid $25 in all. In July, he attended 10 sessions, and he paid $40 in all. How much is the membership fee and how much are the session?

hero (hero):

(5,25) and (10,40) plot and find the slope. Then use the point slope formula

OpenStudy (anonymous):

25=m+5f 40=m+10f -25=-m-5f 15=5f f=3 m=10

OpenStudy (anonymous):

m + 5f = 25 m + 10f = 40 m = 25 - 5f (25 - 5f) + 10f = 40 etc

hero (hero):

or you can do that

OpenStudy (anonymous):

It does say using substitution etc ;)

hero (hero):

I know. I was just showing another way it could be solved.

OpenStudy (anonymous):

(25 - 5f) + 10f = 40 So then I solve this by subracting 5f from both sides right?

OpenStudy (myname):

let x be the total monthly fee let y be the cost per session Then we can get x+5y=25 x=25-5y Again, x+10y=40 Now, we can substitute the value of x and solve for y 25-5y+10y=40 5y=40-25 y=15/5=3 Now , we can find x by using one of the above eq x=25-5(3) x=10 Thus the total monthly fee is 10$ and the fees per session is 3$